Solve the following system of equations for the unknown variables.
a + b = 7
a + c = 12
b + c = 15

Answer by ankor@dixie-net.com(22740)

Solve the following system of equations for the unknown variables.
:
Write it like this:
:
a + b + 0 = 7
a + 0 + c = 12
0 + b + c = 15
:
Multiply the 2nd equation by -1 and add all three:
a + b + 0 = 7
-a -0 - c = -12
0 + b + c = 15
-----------------adding eliminates a and c, find b
0 +2b + 0 = 10
b = 5
:
Very easy to find a and c now